# Soft-Label Caching and Sharpening for Communication-Efficient Federated Distillation

### v1

@@ IV-B Experimental Results
Ablation Study of Cache Duration: The cache duration $D$, a key parameter in our soft-label caching mechanism, significantly influences both communication cost and accuracy in SCARLET. We performed a detailed analysis using CIFAR-10 under a Dirichlet distribution with $\alpha=0.05$, as results with other $\alpha$ values showed similar trends and are therefore omitted for brevity.
The results in Fig. 10 demonstrate that increasing $D$ reduces communication costs for both the server and clients, but it also introduces trade-offs in accuracy. On the server side, $D=800$ makes the system heavily reliant on cached soft-labels (see Fig. 3), causing accuracy degradation. Notably, accuracy shows stepwise improvements at multiples of the cache duration, such as rounds $800$ and $1600$, when the cache is fully refreshed and stale predictions are replaced.
@@
Figure 10: Impact of cache duration $D$ on test accuracy for SCARLET on CIFAR-10 under a Dirichlet distribution with $\alpha=0.05$. Increasing $D$ significantly reduces communication costs for both the server and clients.
Figure 11: Additional analysis of finer cache duration values for SCARLET on CIFAR-10 under a Dirichlet distribution with $\alpha=0.05$. The results confirm $D=200$ as an optimal duration, balancing communication efficiency and accuracy.
On the client side, $D=200$ emerges as an optimal duration, achieving higher accuracy than both shorter and longer durations. As shown in Fig. 11, additional analysis with finer values of $D$ highlights that accuracy consistently improves as $D$ approaches $200$. This balance allows for stable learning periods without frequent soft-label exchanges, enabling efficient utilization of cached predictions and minimizing the risk of overfitting to stale labels.
These findings highlight the trade-off between communication cost savings and accuracy. While longer cache durations like $D=800$ reduce communication costs, they risk accuracy drops due to reliance on outdated cached labels, with improvements only at periodic refresh intervals. Durations around $D=200$ offer a practical balance, enabling stable learning periods without frequent soft-label exchanges and minimizing overfitting to cached predictions. As shown in Fig. 3, such durations can often be identified through simple simulations. In cases where determining an optimal duration is challenging, smaller values of $D$ can still provide substantial communication cost reductions with only modest impacts on accuracy. This flexibility makes cache duration a valuable parameter for tuning communication efficiency in FL systems.
Ablation Study of Enhanced ERA: To evaluate the impact of the aggregation sharpness parameter $\beta$ in the Enhanced ERA mechanism, we conducted experiments on CIFAR-10 under both strong ($\alpha=0.05$) and weak ($\alpha=0.3$) non-IID conditions. A fixed cache duration of $D=50$ was used in all experiments, and the results are presented in Fig. 12.
@@
Figure 12: Impact of aggregation sharpness parameter $\beta$ in Enhanced ERA on test accuracy for SCARLET on CIFAR-10 under Dirichlet distributions with $\alpha=0.05$ (strong non-IID) and $\alpha=0.3$ (weak non-IID). Varying $\beta$ demonstrates distinct effects on server-side and client-side accuracy, highlighting the importance of proper tuning.
@@
For strong non-IID scenarios ($\alpha=0.05$), increasing $\beta$ values (e.g., $\beta>2.0$) significantly improved server-side convergence, reducing the communication costs needed to reach a target accuracy. This acceleration is attributed to the sharper aggregation of client predictions, which mitigates the smoothing effects of locally biased soft-labels. However, on the client side, excessively high $\beta$ values caused accuracy stagnation in later rounds. This degradation likely results from over-sharpened labels introducing biases during local training, thus limiting the generalization of client-specific models.
@@
For weak non-IID conditions ($\alpha=0.3$), $\beta=1.0$ preserved steady server-side accuracy across all communication rounds, avoiding the risks associated with over-sharpening or over-smoothing. On the client side, $\beta=1.5$ yielded the highest accuracy, striking a balance between effective aggregation and model generalization. Larger $\beta$ values led to performance degradation, suggesting that less aggressive sharpening is preferable in scenarios with lower heterogeneity.
These results highlight the necessity of tailoring the value of $\beta$ to the degree of non-IID-ness in client data distributions. For weakly non-IID or near-IID conditions, smaller $\beta$ values (e.g., $\beta\approx 1.0$) are sufficient to maintain stable performance while preventing overfitting. Conversely, in strongly non-IID scenarios, larger $\beta$ values are advantageous as they mitigate the smoothing effects of locally biased soft-labels; however, careful tuning is essential to avoid introducing biases that could negatively impact client model generalization.
@@ V Limitations and Future Directions
First, SCARLET, like other distillation-based FL approaches, depends on the availability of a shared public dataset whose distribution closely resembles, but does not necessarily match exactly, that of the clients’ data. This dependency restricts the applicability scenarios for distillation-based FL methods. Knowledge distillation trains a student model to replicate the outputs of a teacher model for given inputs, essentially enforcing functional equivalence. In theory, even random noise could serve as the shared dataset; however, prior work demonstrates that this approach yields poor results [32]. Recent techniques leveraging generative models have been proposed to produce suitable synthetic data in place of a real public dataset [5], [33], [34]. Nevertheless, training generative models within an FL context still carries a risk of inadvertently exposing patterns of sensitive client data if not carefully managed. Thus, robust safeguards for privacy-preserving synthetic data generation remain critical. Integrating these privacy-conscious data synthesis techniques offers a promising direction to broaden the practical applicability of distillation-based FL frameworks like SCARLET.
@@
Second, the current implementation of SCARLET assumes a homogeneous model architecture across all clients and the server for analytical clarity and simplicity. However, one significant advantage of distillation-based FL is its inherent capability to support heterogeneous models. Several studies have demonstrated successful implementations of model heterogeneity within FL frameworks, covering both in client-server model diversity [35] and inter-client variability [20]. Nevertheless, comprehensive evaluations of accuracy degradation resulting from introducing model heterogeneity remain limited, with only a handful of studies, such as [36] addressing this critical aspect. Consequently, the experimental validation of SCARLET and similar distillation-based FL frameworks under heterogeneous model conditions remains an open research area. Investigating this unexplored dimension is essential not only from an academic standpoint but also for practical FL deployments, where diverse device capabilities necessitate tailored model architectures across client devices.
Third, selecting hyperparameters for the soft-label caching mechanism, particularly the cache duration $D$, entails trade-offs between reduced communication costs and model accuracy. While the current study employs a fixed cache duration, adaptive caching strategies—dynamically adjusting durations based on real-time performance metrics or selectively caching predictions—could further enhance communication efficiency and model performance under diverse FL conditions.
Finally, the Enhanced ERA mechanism presently relies on a static hyperparameter ($\beta$) to regulate aggregation sharpness. Given the inaccessibility of client-side data distributions, future research could explore dynamically adjusting $\beta$ based on aggregated soft-label distributions or observed training dynamics. Such adaptive strategies would improve SCARLET’s robustness and ability to manage varying degrees of data heterogeneity throughout FL.
@@
In summary, future research directions include investigating synthetic data generation methods as alternatives to public datasets, evaluating SCARLET’s efficacy in heterogeneous model settings, and developing adaptive caching strategies responsive to real-time performance metrics. Additionally, refining Enhanced ERA with adaptive aggregation sharpness mechanisms promises further improvements in handling diverse non-IID data scenarios.
@@ VI Conclusion
In this paper, we proposed SCARLET, a novel distillation-based FL framework that addresses key challenges in communication efficiency and data heterogeneity. By introducing the soft-label caching mechanism and the Enhanced ERA mechanism, SCARLET significantly reduces communication costs while maintaining competitive model accuracy. Extensive experiments demonstrated that SCARLET outperforms state-of-the-art distillation-based FL approaches in both server- and client-side evaluations. The modular design of SCARLET ensures compatibility with other FL frameworks, offering a flexible and scalable solution for diverse deployment scenarios. Despite its limitations, SCARLET represents a substantial step forward in communication-efficient FL. By further investigating adaptive caching strategies, model heterogeneity, and dynamic hyperparameter tuning, SCARLET could be extended to address a wider range of real-world challenges. This work lays the groundwork for future research, advancing the development of scalable, efficient, and robust FL frameworks.
@@ Acknowledgments
Appendix A Proof of Majorization for Power Exponents
In this Appendix, we present the mathematical proof of the majorization property introduced in Section III-D. Specifically, we prove that for any two power exponents $\beta_{1}$ and $\beta_{2}$ satisfying $\beta_{2}>\beta_{1}>0$, the distribution produced by Enhanced ERA with exponent $\beta_{2}$ is majorized by the distribution produced with exponent $\beta_{1}$. This result justifies the monotonic reduction of entropy as $\beta$ increases, leveraging the Schur-concavity of Shannon entropy.

### v3

@@ IV-B4 Ablation Study of Cache Duration
The cache duration $D$, a key parameter in our soft-label caching mechanism, significantly influences both communication cost and accuracy in SCARLET. We performed a detailed analysis using CIFAR-10 under a Dirichlet distribution with $\alpha=0.05$, as results with other $\alpha$ values showed similar trends and are therefore omitted for brevity. The results for this ablation study were averaged across three random seeds.
@@
Figure 12: Impact of cache duration $D$ on the accuracy vs. communication cost trade-off (CIFAR-10, $\alpha=0.05$). This figure highlights that while caching ($D>0$) always saves communication, excessively long durations degrade performance. Increasing $D$ from 0 (blue line) to 50 or 100 (green/red lines) significantly shifts the curves left (saving communication) with minimal accuracy loss. However, very long durations (e.g., $D\geq 400$) cause a clear drop in final accuracy, demonstrating the negative impact of using stale cached labels.
The results in Fig. 12 demonstrate the trade-offs associated with cache duration. On the server side (Fig. 12a), we observe that a conservative cache duration provides substantial communication savings with negligible impact on performance. Increasing $D$ from 0 (no cache) up to 200 results in almost no change to the final model accuracy (less than a 1 percentage point difference). However, as the cache duration becomes excessively long (e.g., $D=400$ or $D=800$), the model fails to converge to the target accuracy within the 3000-round training period. This performance degradation is caused by the system’s heavy reliance on stale cached soft-labels, as predicted by our simulation in Section III-C (Fig. 3). Notably, the stepwise accuracy improvements observed for $D=800$ at multiples of its cache duration (e.g., rounds 800 and 1600) support this hypothesis, as these moments correspond to the full refresh of the stale cache.
On the client side (Fig. 12b), a similar trend is observed. For appropriate cache durations (e.g., $D\leq 200$), there is no clear degradation in final accuracy compared to the baseline. However, for excessively long durations (e.g., $D\geq 400$), the training stability markedly decreases, and the model’s performance becomes erratic. Given that these long durations also fail to deliver optimal server-side performance, they are not considered effective.
These findings simplify the practical guidance for tuning $D$. The primary trade-off is not a gradual, nuanced decline, but rather a distinct performance cliff when $D$ becomes excessively large (e.g., $D\geq 400$). This behavior aligns with the risk identified in our earlier simulation (Fig. 3), which predicted that excessively long durations lead to heavy reliance on stale cached soft-labels. The most critical insight for practical deployment is that a conservative duration (e.g., $D=50$ or $D=100$) provides a substantial reduction in communication cost (Fig. 12) with minimal accuracy trade-offs. This demonstrates that a robust and effective $D$ that avoids significant performance loss is easily identifiable, making cache duration a valuable parameter for tuning communication efficiency in FL systems.
IV-B5 Ablation Study of Enhanced ERA
To evaluate the impact of the aggregation sharpness parameter $\beta$ in the Enhanced ERA mechanism, we conducted experiments on CIFAR-10 under both strong ($\alpha=0.05$) and moderate ($\alpha=0.3$) non-IID conditions. A fixed cache duration of $D=50$ was used in all experiments, and the results are presented in Fig. 13.
@@ IV-B5 Ablation Study of Enhanced ERA
Figure 13: Impact of the aggregation sharpness parameter $\beta$ in Enhanced ERA on test accuracy for SCARLET on CIFAR-10 under Dirichlet distributions with $\alpha=0.05$ (strong non-IID) and $\alpha=0.3$ (moderate non-IID). Varying $\beta$ exhibits distinct effects: under strong non-IID conditions (13a, 13b), increasing $\beta$ (e.g., $\beta\geq 1.5$) accelerates server-side convergence and improves accuracy. Under moderate non-IID conditions (13c, 13d), $\beta=1.0$ (simple averaging) is most stable for the server, while $\beta=1.5$ achieves the highest client-side accuracy, highlighting a trade-off between generalization and personalization.
@@
For strong non-IID scenarios ($\alpha=0.05$, Fig. 13a and 13b), increasing $\beta$ values (e.g., $\beta\geq 1.5$) significantly improved server-side convergence (Fig. 13a). While $\beta=1.5$ achieved the highest final accuracy on both server and client, increasing $\beta$ further to 2.0 offered a slight improvement in convergence speed at the cost of a marginal (less than 1 percentage point) drop in final accuracy. This acceleration is attributed to the sharper aggregation of client predictions, which mitigates the smoothing effects of locally biased soft-labels. However, on the client side (Fig. 13b), excessively high $\beta$ values caused accuracy degradation in later rounds. This degradation likely results from over-sharpened labels introducing biases during local training, thus limiting the generalization of client-specific models.
@@
For moderate non-IID conditions ($\alpha=0.3$, Fig. 13c and 13d), $\beta=1.0$ preserved steady server-side accuracy (Fig. 13c) across all communication rounds, avoiding the risks associated with over-sharpening or over-smoothing. On the client side (Fig. 13d), $\beta=1.5$ yielded the highest accuracy, striking a balance between effective aggregation and model generalization. Larger $\beta$ values led to performance degradation, suggesting that less aggressive sharpening is preferable in scenarios with lower heterogeneity.
To more precisely characterize the sensitivity of Enhanced ERA to data heterogeneity, we extend the ablation beyond the two $\alpha$ settings and sweep $\alpha\in\{0.05,\dots,0.30\}$ and $\beta\in\{0.5,\dots,3.0\}$. For each ($\alpha$, $\beta$) pair, we report the mean test accuracy over the last 10 rounds at round 3000 and visualize the results as heatmaps for server- and client-side evaluations (Fig. 14). This grid complements the learning-curve analysis in Fig. 13 and clarifies how sharpening should be scheduled across non-IID strengths.
(a) Server
(b) Client
Figure 14: Sensitivity of Enhanced ERA to data heterogeneity. Heatmaps show final test accuracy (%) vs. aggregation sharpness $\beta$ and Dirichlet $\alpha$ for server-side (14a) and client-side (14b) evaluation. Each cell represents the mean accuracy over the last 10 rounds at round 3000. As $\alpha$ increases (approaching IID), the server-side optimum shifts toward $\beta=1.0$, whereas $\beta=1.5$ remains consistently strong across $\alpha$ and is typically optimal on the client side. Together with Fig. 13, these results highlight a speed–accuracy trade-off: larger $\beta$ accelerates server-side convergence under strong non-IID conditions, while overly sharp aggregation can dampen late-stage client performance gains.
(a) Server
(b) Client
Figure 15: Sensitivity of Enhanced ERA to near-IID data heterogeneity. Heatmaps show final test accuracy (%) vs. aggregation sharpness $\beta$ and Dirichlet $\alpha$ for server-side (15a) and client-side (15b) evaluation. As $\alpha$ increases (approaching IID), the optimal performance for both server and client converges toward $\beta=1.0$, confirming that Enhanced ERA behaves consistently with standard aggregation under homogeneous data conditions.
Three observations emerge consistently. On the server side, the optimal $\beta$ drifts toward 1.0 as the partition approaches IID: under strong non-IID ($\alpha=0.05$), $\beta\approx 1.5\text{-}2.0$ yields the best accuracy, whereas under moderate non-IID ($\alpha=0.3$) the optimum settles at $\beta=1.0$. Notably, $\beta=1.5$ is robust across the entire sweep, remaining within a small margin of the per-$\alpha$ optimum even when that optimum is $\beta=1.0$. On the client side, $\beta=1.5$ consistently attains the highest final accuracy across $\alpha$, and the penalty from “over-sharpening” (increasing $\beta$ by one or two steps) is milder than on the server—particularly when $\alpha$ indicates stronger non-IID.
To provide a complete picture, we extended this evaluation to near-IID scenarios ($\alpha\geq 0.5$), as shown in Fig. 15. This evaluation follows the same methodology, sweeping $\alpha\in\{0.5,\dots,10.0\}$ and $\beta\in\{0.5,\dots,2.0\}$. The results confirm the trend: as the client data distribution approaches IID (i.e., as $\alpha$ increases), the optimal aggregation sharpness $\beta$ converges to 1.0 for both server-side and client-side evaluations.
For $\alpha\geq 1.0$, $\beta=1.0$ (which corresponds to simple soft-label averaging) consistently achieves optimal or near-optimal performance. This indicates that when data heterogeneity is low, the aggressive sharpening ($\beta>1.0$) required for strong non-IID settings is unnecessary and, in fact, degrades performance.
Notably, on the server side (Fig. 15a), for near-IID settings ($\alpha\in\{5.0,10.0\}$), a $\beta$ value below 1.0 (specifically $\beta=0.5$) achieves final accuracy comparable to that of the optimal $\beta=1.0$. This suggests that when all clients produce similar, high-confidence predictions, sharpening provides no additional benefit, and a slight “softening” (or entropy-increasing) effect does not degrade—and may marginally improve—the global model’s accuracy. This implies a potential for Enhanced ERA to be utilized “in reverse” to manage overly confident and homogeneous client predictions, a direction warranting further investigation.
Taken together, these findings (Fig. 13, 14, and 15) demonstrate that Enhanced ERA successfully translates the complex problem of non-IID aggregation into a practical speed-accuracy trade-off. This allows for simple, robust tuning guidelines for deployments where $\alpha$ is unknown: $\beta=1.5$ serves as a strong default, achieving near-optimal for both server and clients across a broad range of heterogeneity. For specific goals, $\beta$ can be adjusted further: prioritizing faster server-side convergence under strong non-IID conditions suggests $\beta=2.0$, while $\beta=1.0$ (i.e., simple averaging) is optimal as the data approaches IID.
IV-B6 Evaluation under Partial Client Participation
We evaluated this partial participation scenario on CIFAR-10 with $K=100$ clients, varying the client participation ratio, denoted as p(0<p<1)p\ (0<p<1), from 0.1 (10 clients per round) to 1.0 (all 100 clients). All other hyperparameters, including $D=50$ and $\beta$ values, remain the same as in the main experiment. The results are summarized in Fig. 16. These graphs plot the final accuracy (mean and standard deviation over the last 10 rounds, left y-axis) and the total cumulative communication cost (right y-axis) against the client participation ratio for both $\alpha=0.05$ (Fig. 16a) and $\alpha=0.3$ (Fig. 16b).
(a) Dirichlet $\alpha=0.05$
(b) Dirichlet $\alpha=0.3$
Figure 16: Performance Comparison for SCARLET under varying client participation ratios ($p$) on CIFAR-10. Solid lines indicate the soft-label caching mechanism was enabled (w/ Cache), while dashed lines indicate it was disabled (w/o Cache). Accuracy values are the mean $\pm$ std. dev. over the final 10 rounds. The figure highlights that enabling caching results in significant communication cost reduction with only slight accuracy degradation in high participation settings ($p\geq 0.7$), whereas in low participation settings ($p\leq 0.3$), it improves accuracy considerably despite increasing communication cost.
As shown in Fig. 16, SCARLET’s caching mechanism (solid lines) demonstrates its primary benefit at high participation ratios ($p\geq 0.7$). It achieves substantial communication cost reductions (e.g., the red solid line is significantly below the red dashed line for $p\geq 0.7$ in both Fig. 16a and Fig. 16b) while maintaining comparable accuracy (blue and green solid lines are close to their dashed counterparts), with performance degradation typically below 1 percentage point in these high participation settings.
Conversely, in low participation scenarios (specifically $p=0.1$), the cache provides a significant stabilization effect. In Fig. 16b, the baseline model’s server accuracy (blue dashed line at $p=0.1$) degrades significantly to 42.51%, whereas the caching mechanism (blue solid line) maintains it at 50.05%, a 7.5 percentage point improvement. We attribute this to temporal stabilization. In low participation settings, the aggregated teacher signal is volatile, as it is computed from a small, frequently changing subset of clients. The baseline model is sensitive to this instability. SCARLET’s cache mitigates this by reusing soft-labels from previous rounds, which functions as a temporal smoothing mechanism, providing a more consistent training signal and dampening the volatility from low client sampling ratios. This stabilization effect introduces a trade-off, as seen in the $\alpha=0.3$, $p=0.1$ scenario (Fig. 16b): the communication cost with the cache (red solid line at 9.99 GB) is significantly higher than without it (red dashed line at 3.36 GB). This increased communication cost is due to the frequent transmission of catch-up packages to stale clients.
These results demonstrate the robustness of SCARLET’s design: while it significantly reduces costs in high participation settings, it also functions as a critical performance stabilizer in low participation, volatile environments.
@@ V Limitations and Future Directions
First, SCARLET, like other distillation-based FL approaches, depends on the availability of a shared public dataset whose distribution closely resembles, but does not necessarily match exactly, that of the clients’ data. This dependency restricts the applicability scenarios for distillation-based FL methods. Knowledge distillation trains a student model to replicate a teacher’s outputs for given inputs, essentially enforcing functional equivalence. In theory, even random noise could serve as the shared dataset; however, prior work indicates that this yields poor results [stanton2021does]. Recent techniques leveraging generative models have been proposed to produce suitable synthetic data in place of a real public dataset [fedgen], [fedcg], [fediod]. Nevertheless, training generative models within an FL context still carries a risk of inadvertently exposing sensitive client patterns if not carefully managed. Thus, robust safeguards for privacy-preserving synthetic data generation remain critical. Integrating these privacy-conscious data synthesis techniques offers a promising direction to broaden the practical applicability of distillation-based FL frameworks like SCARLET.
@@
Second, the current implementation of SCARLET assumes a homogeneous model architecture across all clients and the server for analytical clarity and simplicity. However, one significant advantage of distillation-based FL is its inherent capability to support heterogeneous models. Several studies have demonstrated successful implementations of model heterogeneity within FL frameworks, covering both client-server diversity [fedgkt] and inter-client variability [fedmd]. Nevertheless, comprehensive evaluations of accuracy degradation resulting from introducing model heterogeneity remain limited, with only a handful of studies, such as [cmfd] addressing this critical aspect. Consequently, the experimental validation of SCARLET and similar distillation-based FL frameworks under heterogeneous model conditions remains an open research area. Investigating this unexplored dimension is essential not only from an academic standpoint but also for practical FL deployments, where diverse device capabilities necessitate tailored model architectures across client devices.
Third, our ablation study of the cache duration $D$ (Section IV-B4) clarifies its role and provides strong practical guidance for deployment. The results reveal a distinct performance cliff: conservative $D$ values offer substantial communication savings with minimal accuracy loss, establishing them as a robust and effective default. Conversely, the study confirmed that excessively long durations negatively impact server-side convergence and client-side stability. Critically, this risk is predictable. The lightweight simulation (Section III-C) allows these ineffective, overly long $D$ values to be identified and avoided, before incurring the cost of a full FL experiment. While our current simple and fair refresh strategy proved highly effective with a conservative $D$, a promising future direction is to explore more granular caching policies. For instance, a probabilistic or selective per-sample expiration strategy might mitigate the instability caused by mass-refresh events observed with very long durations, potentially unlocking further efficiency gains.
Finally, our ablation study of Enhanced ERA (Section IV-B5) clarifies its behavior and offers practical guidance for deployment. The mechanism is fundamentally robust, overcoming the class-count ($N$) dependency of conventional ERA and offering stable control where $\beta=1$ recovers simple averaging. Our empirical results also established a clear operational guideline: a static heuristic (e.g., $\beta=1.5$ in our CIFAR-10 tests) achieved near-optimal final accuracy across the entire spectrum of data heterogeneity, from strong non-IID to near-IID. While this establishes a highly effective default, our analysis also revealed opportunities for further optimization, such as prioritizing server-side convergence speed. Therefore, the primary remaining research direction shifts from identifying a static default to automating this tuning process. Developing a theoretical or heuristic method to set $\beta$ adaptively, perhaps using server-visible signals like aggregated soft-label entropy, remains a promising avenue for future work.
@@
In summary, future research directions include investigating synthetic data generation methods as alternatives to public datasets, evaluating SCARLET’s efficacy in heterogeneous model settings, and developing more granular or probabilistic caching policies to build upon our analysis. Additionally, automating the tuning of the Enhanced ERA sharpness parameter $\beta$ promises further improvements by creating a fully dynamic mechanism for diverse non-IID data scenarios.
@@ VI Conclusion
In this paper, we proposed SCARLET, a novel distillation-based FL framework that addresses key challenges in communication efficiency and data heterogeneity. By introducing the soft-label caching mechanism and the Enhanced ERA mechanism, SCARLET significantly reduces communication costs while maintaining competitive model accuracy. Extensive experiments demonstrated that SCARLET outperforms state-of-the-art distillation-based FL approaches in both server- and client-side evaluations. The modular design of SCARLET ensures compatibility with other FL frameworks, offering a flexible and scalable solution for diverse deployment scenarios. SCARLET thus represents a substantial step forward in communication-efficient FL. As discussed in Section V, extending this work by investigating model heterogeneity, developing more granular or probabilistic caching policies, and automating the tuning of the Enhanced ERA sharpness parameter $\beta$ will address an even wider range of real-world challenges.
@@ Acknowledgments
Appendix A Algorithm for Lightweight Cache Hit Rate Simulation
This appendix provides the detailed algorithm for the lightweight, standalone simulation discussed in Section III-C. This simulation is used to estimate the cache hit rate (as shown in Fig. 3) by modeling only the random sampling of the public dataset and the cache expiration logic. As the algorithm demonstrates, it is computationally trivial and runs independently of the FL training, serving as a practical tool for tuning the cache duration $D$.
Algorithm 3 Lightweight Cache Hit Rate Simulation
1:Input:
[…]
4:$D$: Cache duration to simulate (e.g., 50)
5:$T$: Total simulation rounds (e.g., 2000)
6:Output:
[…]
$\triangleright$ Stores the cache timestamp for each sample
[…]
10:if $D=0$ then
11:  $R_{\text{cached}}\leftarrow$ array of size $T$, initialized to 0.0
12:  return $R_{\text{cached}}$
13:end if
14:for $t=1,\cdots,T$ simulation rounds do:
15:  $\mathcal{I}^{t}\leftarrow$ Randomly select $|\mathcal{P}^{t}|$ indices from $1,|\mathcal{P}|]$
16:  $HitCount\leftarrow 0$
17:  for each index $i\in\mathcal{I}^{t}$ do
18:   if $\mathcal{C}_{\text{sim}}[i]=\textbf{null}$ then
19:     $\mathcal{C}_{\text{sim}}[i]\leftarrow t$
      $\triangleright$ Cache miss (empty): Add to cache
20:   else if $t-\mathcal{C}_{\text{sim}}[i]>D$ then
21:     $\mathcal{C}_{\text{sim}}[i]\leftarrow t$
      $\triangleright$ Cache miss (expired): Refresh cache
22:   else
23:     $HitCount\leftarrow HitCount+1$
      $\triangleright$ Cache hit (valid)
24:   end if
25:  end for
26:  Append ($HitCount/|\mathcal{P}^{t}|$) to $R_{\text{cached}}$
27:end for
28:return $R_{\text{cached}}$
Appendix B Proof of Majorization for Power Exponents
In this appendix, we present the mathematical proof of the majorization property introduced in Section III-E. Specifically, we prove that for any two power exponents $\beta_{1}$ and $\beta_{2}$ satisfying $\beta_{2}>\beta_{1}>0$, the distribution produced by Enhanced ERA with exponent $\beta_{2}$ is majorized by the distribution produced with exponent $\beta_{1}$. This result justifies the monotonic reduction of entropy as $\beta$ increases, leveraging the Schur-concavity of Shannon entropy.